OpenStudy (anonymous):

Solve the equation: g/8 = -3/4 -7/8 = 2/3 + n -7t = 5.6 m/8 = 8 1/3 Help thanks:)

OpenStudy (anonymous):

Did you even try any of them?

OpenStudy (anonymous):

Yeah! These are the ones I didn't get I did the other ones.

OpenStudy (anonymous):

ok so for the first one solve for g? what would you do?

OpenStudy (anonymous):

Wouldn't we change - 3/4 to - 4/3?

OpenStudy (anonymous):

\[\frac{ g }{ 8}=\frac{ -3 }{ 4} => g = \frac{ -3 }{ 4 }\times8\]

OpenStudy (anonymous):

g = -6

OpenStudy (anonymous):

why would you change -3/4 to -4/3?

OpenStudy (anonymous):

I thought you have to change the reciprocal

OpenStudy (anonymous):

why would you do that if you're solving for g :S

OpenStudy (anonymous):

Oh.

OpenStudy (anonymous):

you're not moving the -3/4

OpenStudy (anonymous):

you just want to isolate the variables. Now try the second one.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

you want to find n

OpenStudy (anonymous):

right/

OpenStudy (anonymous):

 How do I isloate?